OpenStudy (anonymous):

How many kg of pure salt must be added to 8 kg of a 30% solution to obtain a 35% soltution? Round to the nearest tenth. A. 0.4 B. 0.6 C. 0.8 D. 1.0

OpenStudy (anonymous):

0.3*8 + x = 0.8*0.35 => x = 0.8(0.05) = 0.4